Use the given graphs of x =f(t) and y = g(t) to sketch the corresponding parametric curve in the xy-plane. 1341
Answer #1

Solution: given that The give graphs of a=F(E) - 90 st To 11 selt) = -247 0 t =17 -- 1.4 Wa-t+289 -> from ean o to y = 289-t²y = 289 - (17-x)2 y =289 -289 - x² +34a o y = -x2 +34a hot = x (34 -a) 289/ 34 Nole: Hope this was helpbeel. please do leave